OpenStudy (anonymous):

Which of the following is the simplified form of x minus 7 over 2 x minus 12 plus the fraction 6 over x squared minus 36

OpenStudy (anonymous):

do u mean \[x - 7/2x - 12 + 6 / (x ^{2} - 36)\] ?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

\[\frac{x -7}{2x -12}+\frac{6}{x^2 - 36}\] factorise the denominators \[\frac{ x - 7}{2(x -6)} + \frac{ 6}{(x - 6)(x + 6)}\] now you need to find the common denominator

OpenStudy (anonymous):

is it x + 5/ 2(x + 6)?

OpenStudy (anonymous):

the common denominator will be 2(x-6)(x+6). take LCM.

OpenStudy (campbell_st):

getting a common denominator can be done by the following calculation \[\frac{(x -7)}{2(x -6)} \times \frac{(x +6)}{(x+6} + \frac{6}{(x -6)(x +6)} \times \frac{2}{2}\] which gives \[\frac{(x -7)(x +6)}{2(x-6)(x+6)} + \frac{ 6\times2}{2(x-6)(x+6)}\] next you will need to expand the numerator, collect like terms and cancel any common factors

OpenStudy (campbell_st):

and yes you solution is correct